OpenStudy (anonymous):

What is the value of c so that y = x^2 + 9x + c is a perfect square trinomial? a. 18 b. 9/2 c. 9/4 d. 81/4

OpenStudy (anonymous):

You want y to have the form y = (x+a)^2 where 'a' has to be a number that multiplies together to get 'c', but also adds up to 9. So, we can form a system of linear equations: a * a = c a + a = 9 From the second of those two, one can find that a = 9/2. So, substituting that into the first equation, we have 9/2 * 9/2 = 81/4 = c.

OpenStudy (anonymous):

Thank you so much! I was struggling with this and this is just what I needed!